# A framework for the use of generative modelling in non-equilibrium statistical mechanics

### v4

@@ 3.2 Periodically-firing cells
We will define external $\eta_{t}$ as a stimulus in the world, $s_{t}$ as a sensory state, $\mu_{t}$ as the internal state estimating the phase of the cell, and $a_{t}$ as the state of some actuator or motor acting on the environment to make observations conform to predictions. For simplicty we will assume $s_{t}$ tracks $\eta_{t}$ and that sensor noise is standard white in time, i.e. that the likelihood satisfies
@@
The system’s predictions of sensations given a phase is the cosine function $s(\mu)=A\cos\mu$ and we introduce a predictive control prior making actions prefer to create the encoded preference for the waveform,
@@
The system can treated as a linear Gaussian chain, with Gaussian likelihoods for $s\mid\eta$ and $\eta\mid s,a$, providing an exact expression for the posterior
@@
The reason an inferential interpretation is useful here is in the coding of stimuli in neurones and neuronal ensembles. Minimising variational free energy, and the resultant model of internal and active states behaving as if they were performing Bayesian inference, endows the phase of the oscillation with the interpretation of a posterior belief about expected incoming signals in the environment and the Kalman-like updates as modulations of gain to minimise precision-weighted prediction error, rather than merely a position on a limit cycle. Under the relationship to stochastic thermodynamics forged in e.g. [24], quantitative predictions can be made about the cost of maintaining such beliefs—and with an explicit low-dimensional representation of such signals given by the concentration of the non-equilibrium steady state density, we have an organically emerging neural manifold from those beliefs.
@@ 4 Some preliminary philosophical considerations
We will now enter a more philosophical discussion concerning the use of the FEP, guiding its application and interpretation. The preceding technical arguments establish the FEP as an explanatory principle that can be applied to any thing (particle, person, or population) around which one can draw a boundary or Markov blanket. This means that the FEP is what it says on the tin: it is a principle. Like the principles of least action or of maximum entropy, the FEP is a machinery that produces models [3].
@@
Throughout, ‘as if’ marks an interpretive equivalence: internal flows that minimise variational free energy can be read as variational inference without presupposing explicit, propositional models inside the system. Where specific mechanisms are identified (e.g., precision-weighted message passing), this reading coincides with literal implementation; otherwise, it remains an explanatory stance. How might one apply the FEP to make a model of a particular particle, person or population? The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing. What is observed is simply the blanket states of some ‘thing’ that shield internal states from direct observation. If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed. As noted above, technical term for this relationship between real-world system and generative model is “entailment”: a system that conforms to the dependence structure of the generative model is said to entail that generative model, with non-entailment of a known generative model being surprising.
@@
If we now move to a meta-theoretical perspective and consider one particle (e.g., a scientist or philosopher) observing another particle, we have the interesting situation where the modeller may impute a particular generative model that best explains the mechanics of the observed particle. Is this meta-model a map or a territory? [^4] This meta-theoretical move is commonplace in practical applications of the FEP and is sometimes referred to as meta-Bayesian inference or, more simply, observing the observer [36]. Practically, it involves optimising the parameters of a generative model, such that under the FEP (the ideal Bayesian observer assumption) the observable behaviour of the particle is rendered the most likely. This application of the FEP is often described as computational phenotyping [37]. Using a generative model to capture the formal structure of the environment—and to distinguish it from the model that is carried by particles (the variational density)—can be read as a new take on the nouvelle AI idea that physical systems are their own best models; or as a new take on the good regulator theorem. For instance, in work on robotics following the tradition of embodied cognition [38, 39], practitioners have eschewed the construction of agents with internal representations of some environment. In these systems, the physics and geometry of the situation were sufficient to endow these agents with the capacity to couple to an environment. The world is, on this view, its own best representation. In our view, this approach dovetails nicely with FEP-theoretic modelling, where the generative model in FEP-theoretic constructions is just a joint probability density defined over all the states (or paths) of a system. In other words, the generative model is just our representation of the world as we believe it to be; and this model need not be encoded directly in the particle or agent. Given a generative model or Lagrangian of the appropriate sort, we can show that the system evinces a particular partition (i.e., contains particles); and we can show that subsets of the system track each other, where tracking means inferring or becoming the sufficient statistics of probabilistic beliefs about their external states. We interpret this tracking as a form of inference, namely, variational or approximate Bayesian inference under a generative model (i.e., the map).
@@
In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory. Note here, that the map only exists in relation to the observer—and, crucially, that the observer could be observing itself. In addition, applications of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in. Even if one could, the internal states would be forever sequestered behind blanket states. Applications of the FEP place the generative model centre-stage as explanations for the behaviour of things—in terms of a map of some territory that can only be observed through its impressions on its Markov blanket. Here both observed and observer are described at the mechanistic level by flows on free energy flows; at the informational level, the observer’s higher-order model licenses attributions like ‘the observed system is manipulating an internal model’.
@@ 5 A map of that part of the territory that behaves as if it were a map
Interestingly, because it incorporates this information about things that are coupled and their coupling, it has been suggested that our FEP conflates the metaphorical map (our mathematical model of the thing modelled) and the territory (the thing modelled). The so-called model reification or map-territory fallacy is a general critique of the practice of scientific modelling—in physics, principally due to Cartwright [40, 41]. Some have referenced this fallacy in relation to prior work on the FEP [42, 43]. In particular, it has been suggested that FEP-theoretic modelling conflates the metaphorical “map”—(i.e., the scientific model that scientists use to make sense of some real-world phenomenon)—and “territory,” i.e., the actual physical system that is being modelled. (See [6, 44, 45, 3] for related critical discussions of such claims.) According to this line of thinking, using the FEP to claim that we can model objects as themselves engaging in inference about the statistics of their environment constitutes a case of model reification. The allegation is that, in describing the dynamics of physical objects as implementing a form of inference—as opposed to considering the inferential aspects of the explanation as pertaining to our scientific models of those objects—the FEP theorist mistakenly conflates their metaphorical “map” of the territory (i.e., the scientific, FEP-theoretic model) with the “territory” (the real-world system itself). This constitutes a reification of our model by assuming some aspect of the model or “map” is a real feature of the physical world or “territory” [6, 44]. By clarifying that this potential difficulty does not directly and necessarily apply to FEP-theoretic modelling, we are able to turn the story on its head: namely, we point out that the FEP provides a set of ultimate explanatory constraints on what counts as a model (or map) of any thing that exists in the physical world.
We argue that the FEP can be understood, metaphorically, a “map” of sorts: a map of that part of the territory which behaves or looks as if it were a map. In other words, the FEP provides us with tools to understand the mathematically striking property of self-organising systems: that they look as if they infer, track, reflect, or represent the statistics of their environment. Even more simply, it gives us tools to model systems that look as if they are modelling the world.
@@
The technology of the FEP allows us to say that certain components of the system being considered look as if they track other components, and that this is a feature of our map contained in a generative model under a particular partition. In providing us with a calculus allowing us to model things as modelling other things, the FEP provides us with a map (a scientific model) of any possible map or tracking relationship between things (namely, a mapping between internal and external states, mediated by blanket states). [^5] Here, we are assuming representationalist accounts of scientific practice. For alternatives, see [6].
@@
Does our ability to always describe some dynamics as a form of inference imply that those dynamics are literally a form of inference, as opposed to being a manner of describing those dynamics? Here, it is important to defuse a possible (but in our view naïve) objection. It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories. In developing an inferential account of the dynamics of systems, we need not assume that the particle itself is literally performing inference. [^6] Ways of defending stronger, increasingly literal versions of this positions are available as well. For a defence of the claim that physical systems are quite literally in the business of inference, see [46, 47]. Related to this, views of physical systems implementing computations has been articulated [48, 49], as well as views where systems self-organise by computing predictions [50, 51, 26, 23]. What is at stake with the FEP is an ‘as if’ description. We simply take the fact that surprisal varies with the dynamics of the system, which implies free energy is minimised; the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator.
@@
In summary, our scientific, FEP-theoretic model is, metaphorically, a map that allows us to say that some (internal) subset of the system looks as if it possesses a map; which we interpret formally as tracking the statistics of another (external) subset. Our map (as scientists and modellers) can be identified in an unproblematic way as the generative model of the system, which FEP-theoretic technology enables us to write down. It is by construction that our model is a model of the modelling capacities of subsets of the real-world system: that is, our map is precisely a map of how objects behave as if they were maps, allowing us to construct scientific models of the maps encoded (or looking as if they were encoded) by the internal subset of the system considered. This is an answer to allegations that FEP-theoretic modelling commits the map-territory fallacy: namely, it shows that there is no conflation of the map and the territory, in the sense of the map we postulate as being embodied by the system within the mathematical machinery of the FEP. We remark that we do not claim to invalidate the map-territory fallacy itself. That is, we do not suggest that the map and the territory are never conflated; they obviously are in cases of genuine model reification (as discussed cogently by [44]). Models built using the FEP are still scientific models, and the properties of FEP-theoretic models also run the risk of reification. Instead, we are addressing the claim that the FEP is limited in its effectiveness or engages in model reification because real physical things cannot be modelled as themselves modelling their environments by minimising free energy.
@@
Taking all of the above, we might say that the FEP ultimately provides us with a map of any possible map whatsoever of, or held by, a physical system. This echoes seminal work in philosophy by Wittgenstein [52]. In his famous Tractatus Logico-Philosophicus, Wittgenstein set out to delimit the domain of sensible propositions by determining the general form of a proposition, an utterance able to carry a truth value. This set ultimate constraints on what is sensible, and bounded what is meaningful “from inside” of language itself. Similarly, the FEP provides ultimate constraints on possible maps or modelling relations, which arise from what it means to be a map or model of a physical process at all. Indeed, any modelling process consistent with the laws of physics must conform to the FEP, such that the FEP sets ultimate limits on what can count as a map or model—starting from within the technology of mapping or modelling itself.
@@ 6 Conclusion
We argued that, in describing ‘things’ or ‘particles’ as estimators or (deflated) representations of the systems to which they are coupled, the FEP-theoretic apparatus does not commit the map territory fallacy, i.e., it is false that applications the FEP necessarily reify aspects of the metaphorical map (i.e., our scientific model), mistakenly taking them to be part of the territory (i.e., the real-world system that we wish to model); although this remains possible in principle. We have further argued that this allegation itself constitutes a map-territory fallacy fallacy. In distinguishing the generative model and the variational density—and the generative model and the generative process from the real-world system that we aim to model—the technology of FEP-theoretic modelling allows us to construct a map of that part of the territory which looks to an external observer as if or behaves as if it were a map, without committing model reification. The FEP is thus, metaphorically, a map of any possible map whatsoever of, or held by, a physical system. One ought to celebrate this ‘territory-map mapping,’ and the FEP is ultimately, at its core, a principled approach to the formalisation of this mapping.

### v5

@@ 3.2 Periodically-firing cells
We will define external $\eta_{t}$ as a stimulus in the world, $s_{t}$ as a sensory state, $\mu_{t}$ as the internal state estimating the phase of the cell, and $a_{t}$ as the state of some actuator or motor acting on the environment to make observations conform to predictions. For simplicity we will assume $s_{t}$ tracks $\eta_{t}$ and that sensor noise is standard white in time, i.e. that the likelihood satisfies
@@
The system’s predictions of sensations given a phase follow the cosine function $s(\mu)=A\cos\mu$ and we introduce a predictive control prior making actions prefer to create the encoded preference for the waveform,
@@
The system can be treated as a linear Gaussian chain, with Gaussian likelihoods for $s\mid\eta$ and $\eta\mid s,a$, providing an exact expression for the posterior
@@
The reason an inferential interpretation is useful here is in the coding of stimuli in neurones and neuronal ensembles. Minimising variational free energy, and the resultant model of internal and active states behaving as if they were performing Bayesian inference, endows the phase of the oscillation with the interpretation of a posterior belief about expected incoming signals in the environment and the Kalman-like updates as modulations of gain to minimise precision-weighted prediction error, rather than merely a position on a limit cycle. Under the relationship to stochastic thermodynamics forged in e.g. [parr2020markov], quantitative predictions can be made about the cost of maintaining such beliefs—and with an explicit low-dimensional representation of such signals given by the concentration of the non-equilibrium steady state density, we have an organically emerging neural manifold from those beliefs.
3.3 Looking ahead
It is important to consider that we have not quite taken an agnostic approach, i.e., taken an existing model, extracted a free energy gradient from it, and then tested our principle on it. We begin with a generative model and a choice of non-equilibrium steady state, and then engineer a Markov-blanketted Langevin dynamics of the form (2.1.1) with prescribed $Q$ and $\Gamma$, showing in detail how these dynamics can be written as a flow on variational free energy. Since the periodically firing cells model above has been crafted in a bespoke fashion to have a prescribed $Q$, $\Gamma$, and non-equilibrium steady state, it would be interesting to contrast this with a given set of equations, validated by independent modelling work, and treat it using the above reasoning. That is to say, the example was explicitly designed so that the solenoidal term $Q$ and dissipative term $\Gamma$ are convenient (and, for simplicity, taken to be constant), making the inferential structure and thermodynamic quantities analytically transparent. This means that the example is not an independent empirical test of the FEP, but a carefully crafted illustration of what the formalism looks like when the required decomposition is assumed to exist. However—this top-down approach illustrates how our FEP might be used to model a system and explain some observed behaviours inferentially, once a dynamical system built around those behaviours is given. The examples in §3.1–3.2 are explicit constructions that exploit this fact: starting from a chosen non-equilibrium steady state and generative model, they show how the FEP provides a tractable representation of self-organisation (to a fixed point or limit cycle) as approximate Bayesian inference. A more demanding, ‘bottom-up’ demonstration would begin from an independently specified stochastic dynamical model—for example, a physically motivated Langevin equation—and then: (i) derive its steady-state density $p^{*}(x)$ and probability current, (ii) determine whether there exist fields $Q(x)$ and $\Gamma(x)$ satisfying the constraints of (2.1.1), and if so, (iii) read off the corresponding free energy functional and variational interpretation. This is precisely the kind of agnostic derivation that would address worries about circularity: instead of constructing a system to fit a free-energy-minimising flow, one would show that a pre-existing model can be recast in this way under clearly stated assumptions.
Our central contribution in the present paper is therefore more modest but, we would argue, still substantive. Namely, we show that, whenever a decomposition such as (2.1.1) and a Markov blanket exist, one can construct an associated family of generative models and variational densities such that the physical dynamics can be written as descending a variational free energy functional. On this basis, the two toy systems are offered as worked examples of this constructive use of the FEP, rather than as empirical validations of its universality.
@@ 4 Some preliminary philosophical considerations
We will now enter a more philosophical discussion concerning the use of the FEP, guiding its application and interpretation. The preceding technical arguments establish the FEP as an explanatory principle that can be applied to any thing (particle, person, or population) around which one can draw a boundary or Markov blanket. This means that the FEP is what it says on the tin: it is a principle. Like the principles of least action or of maximum entropy, the FEP is a piece of machinery that produces models [ramstead2022bayesian].
@@
Throughout, ‘as if’ marks an interpretive equivalence: internal flows that minimise variational free energy can be read as variational inference without presupposing explicit, propositional models inside the system. Where specific mechanisms are identified (e.g., precision-weighted message passing), this reading coincides with literal implementation; otherwise, it remains an explanatory stance. How might one apply the FEP to make a model of a particular particle, person or population? The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing. What is observed is simply the blanket states of some ‘thing’ that shield internal states from direct observation. If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed. As noted above, the technical term for this relationship between real-world system and generative model is “entailment”: a system that conforms to the dependence structure of the generative model is said to entail that generative model, with non-entailment of a known generative model being surprising.
@@
If we now move to a meta-theoretical perspective and consider one particle (e.g., a scientist or philosopher) observing another particle, we have the interesting situation where the modeller may impute a particular generative model that best explains the mechanics of the observed particle. Is this meta-model a map or a territory? [^4] This meta-theoretical move is commonplace in practical applications of the FEP and is sometimes referred to as meta-Bayesian inference or, more simply, observing the observer [daunizeau]. Practically, it involves optimising the parameters of a generative model, such that under the FEP (the ideal Bayesian observer assumption) the observable behaviour of the particle is rendered the most likely. This application of the FEP is often described as computational phenotyping [schwartenbeck]. Using a generative model to capture the formal structure of the environment—and to distinguish it from the model that is carried by particles (the variational density)—can be read as a new take on the nouvelle AI idea that physical systems are their own best models; or as a new take on the good regulator theorem. For instance, in work on robotics following the tradition of embodied cognition [beer1995computational, brooks1991intelligence], practitioners have eschewed the construction of agents with internal representations of some environment. In these systems, the physics and geometry of the situation were sufficient to endow these agents with the capacity to couple to an environment. The world is, on this view, its own best representation. In our view, this approach dovetails nicely with FEP-theoretic modelling, where the generative model in FEP-theoretic constructions is just a joint probability density defined over all the states (or paths) of a system. In other words, the generative model is just our representation of the world as we believe it to be; and this model need not be encoded directly in the particle or agent. Given a generative model or Lagrangian of the appropriate sort, we can show that the system evinces a particular partition (i.e., contains particles); and we can show that subsets of the system track each other, where tracking means inferring or becoming the sufficient statistics of probabilistic beliefs about their external states. We interpret this tracking as a form of inference, namely, variational or approximate Bayesian inference under a generative model (i.e., the map).
@@
In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory. Note here, that the map only exists in relation to the observer—and, crucially, that the observer could be observing itself. In addition, the application of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in. Even if one could, the internal states would be forever sequestered behind blanket states. Applications of the FEP place the generative model centre-stage as explanations for the behaviour of things—in terms of a map of some territory that can only be observed through its impressions on its Markov blanket. Here both observed and observer are described at the mechanistic level by flows on free energy gradients; at the informational level, the observer’s higher-order model licenses attributions like ‘the observed system is manipulating an internal model’.
@@ 5 A map of that part of the territory that behaves as if it were a map
Interestingly, because it incorporates this information about things that are coupled and their coupling, it has been suggested that our FEP conflates the metaphorical map (our mathematical model of the thing modelled) and the territory (the thing modelled). The so-called model reification or map-territory fallacy is a general critique of the practice of scientific modelling—in physics, principally due to Cartwright [cartwright1984laws, cartwright1989nature]. Some have referenced this fallacy in relation to prior work on the FEP [Bruineberg2020, VanEs2020]. In particular, it has been suggested that FEP-theoretic modelling conflates the metaphorical “map”—(i.e., the scientific model that scientists use to make sense of some real-world phenomenon)—and “territory,” i.e., the actual physical system that is being modelled. (See [Andrews2021, andrews2022making, kirchhoff2022literalist, ramstead2022bayesian] for related critical discussions of such claims.) According to this line of thinking, using the FEP to claim that we can model objects as themselves engaging in inference about the statistics of their environment constitutes a case of model reification. The allegation is that, in describing the dynamics of physical objects as implementing a form of inference—as opposed to considering the inferential aspects of the explanation as pertaining to our scientific models of those objects—the FEP theorist mistakenly conflates their metaphorical “map” of the territory (i.e., the scientific, FEP-theoretic model) with the “territory” (the real-world system itself). This constitutes a reification of our model by assuming some aspect of the model or “map” is a real feature of the physical world or “territory” [Andrews2021, andrews2022making]. By clarifying that this potential difficulty does not directly and necessarily apply to FEP-theoretic modelling, we are able to turn the story on its head: namely, we point out that the FEP provides a set of ultimate explanatory constraints on what counts as a model (or map) of any thing that exists in the physical world.
We argue that the FEP can be understood, metaphorically, as a “map” of sorts: a map of that part of the territory which behaves or looks as if it were a map. In other words, the FEP provides us with tools to understand the mathematically striking property of self-organising systems: that they look as if they infer, track, reflect, or represent the statistics of their environment. Even more simply, it gives us tools to model systems that look as if they are modelling the world.
@@
The technology of the FEP allows us to say that certain components of the system being considered look as if they track other components, and that this is a feature of our map contained in a generative model under a particular partition. In providing us with a calculus allowing us to model things as modelling other things, the FEP provides us with a map (a scientific model) of any possible map or tracking relationship between things (namely, a mapping between internal and external states, mediated by blanket states). [^5] Here, we are assuming representationalist accounts of scientific practice. For alternatives, see [Andrews2021].
@@
Does our ability to always describe some dynamics as a form of inference imply that those dynamics are literally a form of inference, as opposed to being a manner of describing those dynamics? Here, it is important to defuse a possible (but in our view naïve) objection. It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories. In developing an inferential account of the dynamics of systems, we need not assume that the particle itself is literally performing inference. [^6] Ways of defending stronger, increasingly literal versions of this positions are available as well. For a defence of the claim that physical systems are quite literally in the business of inference, see [Kiefer2017literal, Kiefer2020psycho]. Related to this, views of physical systems implementing computations has been articulated [horsman2014does, aaronson2005guest], as well as views where systems self-organise by computing predictions [still2012thermodynamics, perunov2016statistical, ueltzhoffer2021drive, seifert2019stochastic]. What is at stake with the FEP is an ‘as if’ description. We simply take the fact that surprisal varies with the dynamics of the system, which implies free energy is minimised; the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator.
@@
In summary, our scientific, FEP-theoretic model is, metaphorically, a map that allows us to say that some (internal) subset of the system looks as if it possesses a map; which we interpret formally as tracking the statistics of another (external) subset. Our map (as scientists and modellers) can be identified in an unproblematic way as the generative model of the system, which FEP-theoretic technology enables us to write down. It is by construction that our model is a model of the modelling capacities of subsets of the real-world system: that is, our map is precisely a map of how objects behave as if they were maps, allowing us to construct scientific models of the maps encoded (or looking as if they were encoded) by the internal subset of the system considered. This is an answer to allegations that FEP-theoretic modelling commits the map-territory fallacy: namely, it shows that there is no conflation of the map and the territory, in the sense of the map we postulate as being embodied by the system within the mathematical machinery of the FEP. We remark that we do not claim to invalidate the map-territory fallacy itself. That is, we do not suggest that the map and the territory are never conflated; they obviously are in cases of genuine model reification (as discussed cogently by [andrews2022making]). Models built using the FEP are still scientific models, and the properties of FEP-theoretic models also run the risk of reification. Instead, we are addressing the claim that the FEP is limited in its effectiveness or engages in model reification because real physical things cannot be modelled as themselves modelling their environments by minimising free energy.
@@
Taking all of the above, we might say that the FEP ultimately provides us with a map of any possible map whatsoever of, or held by, a physical system. This echoes seminal work in philosophy by Wittgenstein [wittgenstein2013tractatus]. In his famous Tractatus Logico-Philosophicus, Wittgenstein set out to delimit the domain of sensible propositions by determining the general form of a proposition, an utterance able to carry a truth value. This set ultimate constraints on what is sensible, and bounded what is meaningful “from inside” of language itself. Similarly, the FEP provides ultimate constraints on possible maps or modelling relations, which arise from what it means to be a map or model of a physical process at all. Indeed, any modelling process consistent with the laws of physics must conform to the FEP, such that the FEP sets ultimate limits on what can count as a map or model—starting from within the technology of mapping or modelling itself.
@@ 6 Conclusion
We argued that, in describing ‘things’ or ‘particles’ as estimators or (deflated) representations of the systems to which they are coupled, the FEP-theoretic apparatus does not commit the map territory fallacy, i.e., it is false that applications of the FEP necessarily reify aspects of the metaphorical map (i.e., our scientific model), mistakenly taking them to be part of the territory (i.e., the real-world system that we wish to model); although this remains possible in principle. We have further argued that this allegation itself constitutes a map-territory fallacy fallacy. In distinguishing the generative model and the variational density—and the generative model and the generative process from the real-world system that we aim to model—the technology of FEP-theoretic modelling allows us to construct a map of that part of the territory which looks to an external observer as if or behaves as if it were a map, without committing model reification. The FEP is thus, metaphorically, a map of any possible map whatsoever of, or held by, a physical system. One ought to celebrate this ‘territory-map mapping,’ and the FEP is ultimately, at its core, a principled approach to the formalisation of this mapping.